n 110. If n, p, and x are positive integers, and 2. +2, P. then P n 1 E. x + 2 2 x + 2 F. G. х 2x+2 1 H. + 2
Answer #1

Answer:

Given expression:

\frac{n}{p} = \frac{2}{x} + 2 where n, p, and x are positive integers.

\frac{p}{n} = ?

Using the above information, we can find the required value as -

\frac{n}{p} = \frac{2}{x} + 2

\frac{n}{p} = \frac{2 +2x}{x}

\therefore \ \frac{p}{n} = \frac{x}{2 +2x}

Hence, the correct option is (G) .